OpenStudy (anonymous):

consider the following quadratic equation: 10x^2+19x+9=3 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side of the equation into two linear factors.

OpenStudy (anonymous):

10*(x+1)*(x+.9) = 3

OpenStudy (anonymous):

using the standard form ax^2+bx+c=0

OpenStudy (anonymous):

10*(x+0.4)*(x+1.5) = 0